OpenStudy (anonymous):

I want to build a right triangular garden on the side of my house. Find the sides of the triangle if the hypotenuse is 6 feet and the two sides are equal in length

OpenStudy (anonymous):

a^2+b^2 = c^2 a=b so.. a^2+a^2=c^2 c^2 = 6x6 =36 Simplified that's: 2a^2=36 divide both sides by 2 a^2=18 Then get the square root of both sides a & b = 4.24

OpenStudy (anonymous):

thank you so much I get it now

OpenStudy (anonymous):

You Are Very Welcome (:

OpenStudy (anonymous):

Convert 5 times pi divided by 6 radians into degrees which formula do you use? degrees=radians 180/n or the other one